### v1

@@ 4.4 Vortex dynamics and phase slips in toroidal BEC
The persistent flow is indeed started by quantum phase slips of the superfluid order parameter $\psi(r)=\sqrt{\rho(r)}e^{i\phi(r)}$ occurring in the proximity of the region where the superfluid is stirred and accompanied by production and motion of vortices. Quantum phase slips and vortices determine quantized winding numbers of the superfluid flow and its decay[337, 366, 363, 355, 361, 360]. Vortices are generated as a result of a non-uniform backflow along the cross-section of the torus; the backflow itself occurs as the superfluid tends to fill the matter’s low-density region created by the stirring [370]. A semi-classical picture, obtained by equating the velocity of the flow and the critical sound speed inside the barrier, gives an accurate estimate of the conditions for vortex excitations to enter the ring [392]. By increasing the height of the potential barrier set by the stirring laser, the backflow can be reduced, affecting in turn vortex and phase slip dynamics. The quantitative study of such barrier height dependence was carried out by Yakimenko’s group in a series of works based on 3D and 2D GPE [393, 394, 395] –see Fig. 16. The analysis indicates that, for moderate barrier heights, a vortex-antivortex pair is in the immediate spatial position of the stirring laser beam; then, above a critical angular velocity of the stirring, the vortex-antivortex pair unbinds, with the vortices propagating in the condensate at nearly the speed of sound; finally, the antivortex decays as result of dissipation and the vortex is effectively pinned at the center of the ring and sets the persistent flow. For a stronger barrier, the vortex-antivortex pair is formed again close to the stirring potential, but the whole pair moves around the internal side of the torus, rotating in the opposite direction to the stirring; in this rotation, vortices transfer their winding number to the condensate that is eventually set into motion.
@@
 [^13] The effects caused by a red-detuned attractive stirring potential were also analysed [396]. In this case, theoretical simulations predict that the superfluid is characterized by a specific phononic decay with suppression of vortex-antivortex pairs nucleation [397].
Phase slip dynamics was demonstrated to provide the underlining mechanism to set a persistent current also in non-ring geometries as racetracks confinements [398]. For stirring barrier strengths above a critical value, transitions from curved to straight racetrack parts are characterized by a circulation that results in to increase of one unit; in the transitions from straight to curved parts, instead the circulation results in to decrease of one unit. The analysis theoretically demonstrated that, playing with stirring barrier height, stir speed and racetrack geometry, quantized smooth atomic flows can be engineered on–demand in such systems. The interplay between curvature and atom-atom interactions in elliptic systems was explored in Refs. [399, 400].
@@
As was discussed in Sec. 4.1, the persistent flow is characterized by a hysteretic behavior as shown Ref. [360] and shown in Fig. 12: when a rotating flow is made to decrease, the transition to lower winding numbers occurs at stirring frequencies smaller than the ones that were imparted to achieve the original rotation one started from [360]. The hysteresis results in to decrease by increasing the barrier strength. While the quantitative understanding of such a phenomenon is still missing, there is a consensus in the community that hysteresis occurs as a result of the irreversibility encoded in vortex decay [368, 395, 401]. In Ref. [360], hysteresis among different branches is observed and proposed as a method to control an atomtronics device. In the mean-field regime, one can also investigate how hysteresis can be understood when following the ground state solution of a stirred BEC. In particular, in [401, 402], an exact analysis through GPE was shown to produce unstable solutions forming loops in the energy landscape of the system. As shown in Fig. 17, this leads to the facilitation of phase slips to access more stable states as the velocity of the barrier is increased above a critical velocity.
Motivated by the implementations of the atomic counterpart of the dc-SQUID [132, 133], two moving weak links on the ring in opposite directions have been studied [363]. For such a configuration, vortex formation and phase slip dynamics explaining the current-phase relation was studied [397, 134]. Bosonic ring condensates with radii of a few hundred microns have been considered, therefore demonstrating the remarkable progress achieved in the coherent manipulation of condensates that can be seen with the naked eye.
@@ 5.1 BCS-BEC cross-over
A mixture of weakly attractive two-component Fermi gases forms pairs at opposite momenta around the Fermi sphere. This is the renowned Bardeen-Cooper-Schrieffer (BCS) state, which has been introduced to explain superconductivity. By tuning to stronger values the attraction among the fermions, the Cooper pairs modify onto smaller and finally close to point-like bosons, which can undergo to Bose-Einstein condensation of such composite bosons. This is the celebrated BCS-BEC crossover, which plays an important role in different domains, ranging from nuclear to condensed matter physics. Ultracold atomic gases provide a unique platform to explore the BCS-BEC crossover as the interactions are tunable. This allows in particular to test different theoretical approximations. In the BCS regime, the correlation length of the bound pairs is large compared with the typical interparticle distance; in the BEC regime, instead, the pairs are tightly bound in the real space and the pair correlation length is much smaller than the distance between the particles. At intermediate values of the BCS-BEC crossover, one reaches a regime of infinitely attractive interactions, or unitary limit, where the gas properties are fixed just by the density and temperature of the gas.
@@ 5.2 Experimental results
Figure 18: (a) Absorption image of an equal spin mixture of lithium atoms. (b) Experimental radial column density. The black dotted line is the expected density profile for an ideal Fermi gas. In region I, the system is 3D degenerate, II is quasi-2D degenerate, and III is quasi-2D thermal. Figure from Ref. [47].
@@
Faster interaction-driven ramps across the normal-to-superfluid phase transition were observed to result in spontaneous current formation, even for a non-rotating normal ring. This motivated the Dartmouth group to measure the rate of spontaneous current formation as a function of the quench rate, and compare the results against predictions obtained from Kibble-Zurek theory as shown in Ref. [335] and Fig. 19. In these experiments they drove thermal quenches as fast as 50 ms without substantial change in the peak density and chemical potential of the ring by using a hybrid quench protocol where the interaction strength and trap depth were ramped simultaneously. The trap geometry was a concentric double-ring, with one ring treated as a reference ring for interferometric detection of the current state in the other ring. For the fastest quench rates, the freeze-out time was reached before beyond-mean-field effects become significant and the spontaneous current formation rate matched well with predictions of mean-field Kibble-Zurek theory. For slower quenches, the rate of current production declined more rapidly. This was attributed to finite size effects, with the correlation length becoming comparable to the ring circumference for sufficiently slow quench rates.
@@
Figure 22: (a) After preparing an initial persistent current state with circulation $w_{0}$, the $n$ junctions are ramped up. The 3D density plots are isosurfaces obtained from 3D GPE numerical simulations of the experimental setup. If $n<n_{c}$ is below a critical value, the initial current is dissipated via the nucleation of vortices. Conversely, if $n\geq n_{c}$, the system remains stable with $w=w_{0}$ (lower right plot). (b) Examples of single-shot experimental in-situ images and interferograms obtained for $w_{0}=2$ and for the same number of junctions $n$ as in (a): n = 2 (unstable configuration), at $t=0$ (i), $t=1$ms (ii) and $t=7$ms (iii); and $n=4$ (stable configuration) for $t=0$ms (iv), $t=1$ms (v) and $t=20$ms (vi). In the case (iii), the circulation has decayed ($w(t)<w_{0}$) and the vortex emission is identified by the single spiral arm and the presence of a vortex. Figure from Ref. [239].
@@ 5.3 Theoretical Understanding
Still in its infancy, the theoretical study of persistent currents of ultracold matter is currently restricted to one-dimensional rings. At zero temperature and in the absence of interactions, the total energy of a spinless fermionic system is given by
@@
with $k_{n}=\frac{2\pi}{L}(n-\phi/\phi_{0})$ being the associated momentum of the eigenstates and $\{n\}$ being the set of quantum numbers corresponding to the energy levels occupied by the particles as shown in Fig. 23141414The set of quantum numbers $\{n\}$ is inherently related to the charge quantum number $I_{j}$ parameterizing the Bethe equations with $I_{j}=n$ or $I_{j}=n+\frac{1}{2}$ with $j=1,\ldots,N_{p}$. By construction, the sum of these charge quantum numbers is integer and is related to the angular momentum per particle as $\ell=N\frac{\sum_{j}I_{j}}{N_{p}}$.. In contrast with bosons, only $N$ fermions can occupy a single-particle energy level due to the particle statistics. The resulting particle distribution, centered compactly around zero to minimize the energy, is symmetric or asymmetric depending on whether the number of particles is $N_{p}=(2m+1)N$ or $N_{p}=(2m)N$ for integer $m$ giving rise to the parity effect: a diamagnetic or paramagnetic current respectively –Fig. 23. Such a phenomenon is the straightforward generalization of the facets of the Leggett theorem discussed in Sec. 3.3 to $N$-component fermions [161, 85]. As the effective magnetic flux threading the system increases, the Fermi sphere is displaced so as to counteract it and minimize the energy as per Eq. (97[164, 87].
Figure 23: Schematic representation of the energy level occupations for non-interacting $N_{p}$ SU(2) fermions and their behaviour with increasing angular momentum per particle $\ell$. To counteract the increase in $\ell$, the Fermi sphere displacement acquires a shift. Naturally, due to the Pauli exclusion, such a displacement for larger $N_{p}$ is mediated by going to higher $\ell$. Image reprinted from [87].
@@
Figure 24: Interference dynamics for weakly interacting SU($N$) fermions. (a) Homodyne protocol: Momentum distribution $\sum_{\{n\}}|J_{n}(\mathbf{k})|^{2}$ re-scaled by the number of components $N$. At angular momentum per particle $\ell=0$, the momentum distribution is peaked at the origin. Nevertheless, for $\ell>0$ there is a collapse, and the characteristic hole appears. Self-heterodyne protocol: Ring-centre interference $G_{R,C}$ for $N_{p}=6$ particles as a function of $N$ and $\ell$ at short time expansion $t=0.025$. Just as in its homodyne counterpart, the appearance of spirals necessitates that the Fermi sphere is displaced by half. For a fixed number of particles ($N_{p}=6$ in this case), they emerge at lower $\ell$. Additionally, dislocations (radially segmenting white lines) characterize the interferograms of different $N$. The colour bar is non-linear by setting $\mathrm{sgn}(G_{R,C})\sqrt{|G_{R,C}|}$. Images taken from [87].
@@
Figure 25: Persistent current $I(\phi)$ of SU($N$) symmetric fermions with repulsive interactions $U$ as a function of the synthetic gauge field $\phi$ for different filling fractions. (a) Incommensurate regime: Current exhibits a reduced periodicity when going to strong $U$. The quantities $\phi_{d}$ and $\phi_{s}$ correspond to the degeneracy point and flux at which spinons are generated respectively. (b) Mott regime: Persistent current as a function of $U$. Changes in its profile (magnitude and shape) capture the onset to the Mott phase transition. Insets in both figures depict the Bethe ansatz energies $E(\phi/\phi_{0})$ of the Gaudin-Yang-Sutherland and Lai-Sutherland model for various sets of quantum numbers $X$. Figure taken from [85].
@@
The qualitative and quantitative changes of the persistent current make it a suitable probe to distinguish between the various and to map out the system’s phase diagram [85, 89]. The current exhibits critical behaviour during a quantum phase transition despite its mesoscopic nature. For the case in question, the onset to the Mott phase transition in SU($N>2$) Hubbard models was demonstrated to be captured with the persistent current through finite-size scaling analysis [85]. Studies of Mott phase transitions with currents were also extended to multi-orbital SU($N$) Hubbard models to construct a conductivity-based phase diagram [93, 94, 92]. In contrast with the single-band model, it was shown that the current’s suppression with interaction is not always the case. Instead, it thrives in the deep insulating state, and in certain regions known as Hund’s metal region, it is also enhanced as the hopping processes across the different orbitals facilitate the fermions’ motion even at strong interactions.
@@
Figure 28: Homodyne protocol for strongly interacting SU($N$) fermionic systems. (a) Momentum distribution cross-section $n(k_{x},0)$ for large repulsive interactions $U=10,000$ as a function of the effective flux $\phi$. The initially peaked momentum distribution collapses at $k_{x}=0$ once the additional threshold constraint $(N_{p}-1)/2N_{p}$ enforced by the fractionalization is cleared. Insets correspond to the full momentum distribution. (b) Variance of the width of the momentum distribution $\sigma_{n_{k}}(\phi)$ for strongly attractive fermions. The appearance of discrete steps per flux quantum $\phi_{0}$ equal to the number of components $N$ corresponds to the number of particles composing the bound state and, in turn, reflects the angular momentum fractionalization. Figure taken from [87].
In the presence of finite temperature, changes in the persistent current periodicity stem from the interplay between interactions and thermal fluctuations. Employing a thermodynamic Bethe ansatz approach for repulsively interacting two-component fermions, it was demonstrated that apart from having its sawtooth shape smoothened out, the persistent current’s amplitude is also exponentially reduced with the decay rate dependent on the system’s parameters [410]. The suppression of the current could be attributed to temperature activated transitions between states with different angular momenta. Alternatively, the persistent current contribution from the populated levels are equal in magnitude but with different directionalities thereby cancelling each other out. A remarkable effect of finite temperature is that the single-particle frequency of the persistent current is re-instated, imitating, in a way, a reduction in the interaction strength – Fig. 27(b). This ‘de-fractionalization’ phenomenon was also demonstrated to hold for $N$-component fermions with attractive interactions [86]. The mechanism behind this is that the contributing excited states have different frequencies that modify the current’s periodicity, with the general trend being that the system becomes more robust to thermal fluctuations with increasing interactions. Nevertheless, for strong attractive interactions the system is more susceptible to temperature. It turns out that the bound states sub-band, which is the only contributor to the current in this regime, flattens out due to a decrease in the level spacing between the states constituting it. Consequently, at small temperatures the system exists as a ‘gas of bound states’ altering the persistent current periodicity in the same manner discussed previously.
@@
To round things off, we consider the read-out of persistent currents in the strongly interacting regimes through interferometric means [^15] From the theoretical standpoint, the calculation of interference patterns for strongly interacting systems is limited to small particle numbers and system sizes. Unlike the free fermion regimes, there is no exact expression for the correlators in this regime and these systems are intractable by numerical methods such as exact diagonalization and DMRG for large particle numbers and system sizes. Nevertheless for the strongly repulsive interacting regime, it was shown that by employing a Bethe ansatz scheme to split the SU($N$) Hubbard model into the spinless fermionic and SU($N$) Heisenberg models, interference patterns for systems well beyond the current state-of-the-art numerical methods can be obtained [88].. In the homodyne protocol, the fractionalization in the system generates a momentum distribution depression, i.e. a non-zero local minimum in the origin, instead of the characteristic hole. Furthermore, the appearance of this depression occurs at larger flux values compared to the zero and weak interaction picture. On top of the Fermi sphere effect explained previously, there is an additional delay of $\frac{N_{p}-1}{2N_{p}}$ ($\frac{N-1}{2N}$) in strongly repulsive (attractive) regimes, which is a signature of the fractionalization –Fig. 28(a). The visibility of the depression for attracting fermions is significantly diminished due to the reduced coherence associated with the bound state formation. Nevertheless, information can be acquired indirectly from the variance of the width of the momentum distribution, with the appearance of discrete steps whose number in a given flux quantum correspond to the angular momentum fractionalization –Fig. 28(b). For the self-heterodyne protocol, the interferograms also embody the features of the fractionalized with the emergence of the spiral corresponding with the depression in the momentum distribution. Aside from being able to track the angular momentum quantization by spiral’s specific delay, it can also be monitored through the change in the number and orientation of the dislocations. At variance with free fermions, the dislocations in the strong interaction regime are flux dependent being modified by the presence of spin excitations giving a visual cue about the angular momentum fractionalization –Fig. 29. In the case of attractive interactions, self-heterodyne interference patterns are not a good figure of merit for SU($N>2$) systems since the features of $N$-body bound are not captured by the density-density correlators used to generate these interferograms.

### v3

@@ 4.4 Vortex dynamics and phase slips in toroidal BEC
The persistent flow is indeed started by quantum phase slips of the superfluid order parameter $\psi(r)=\sqrt{\rho(r)}e^{i\phi(r)}$ occurring in the proximity of the region where the superfluid is stirred and accompanied by production and motion of vortices. Quantum phase slips and vortices determine quantized winding numbers of the superfluid flow and its decay[362, 390, 387, 379, 385, 384]. Vortices are generated as a result of a non-uniform backflow along the cross-section of the torus; the backflow itself occurs as the superfluid tends to fill the matter’s low-density region created by the stirring [394]. A semi-classical picture, obtained by equating the velocity of the flow and the critical sound speed inside the barrier, gives an accurate estimate of the conditions for vortex excitations to enter the ring [420]. By increasing the height of the potential barrier set by the stirring laser, the backflow can be reduced, affecting in turn vortex and phase slip dynamics. The quantitative study of such barrier height dependence was carried out by Yakimenko’s group in a series of works based on 3D and 2D GPE [421, 422, 423] –see Fig. 16. The analysis indicates that, for moderate barrier heights, a vortex-antivortex pair is in the immediate spatial position of the stirring laser beam; then, above a critical angular velocity of the stirring, the vortex-antivortex pair unbinds, with the vortices propagating in the condensate at nearly the speed of sound; finally, the antivortex decays as result of dissipation and the vortex is effectively pinned at the center of the ring and sets the persistent flow. For a stronger barrier, the vortex-antivortex pair is formed again close to the stirring potential, but the whole pair moves around the internal side of the torus, rotating in the opposite direction to the stirring; in this rotation, vortices transfer their winding number to the condensate that is eventually set into motion.(11)(11)(11)The effects caused by a red-detuned attractive stirring potential were also analysed [424]. In this case, theoretical simulations predict that the superfluid is characterized by a specific phononic decay with suppression of vortex-antivortex pairs nucleation [425].
@@
Phase slip dynamics was demonstrated to provide the underlining mechanism to set a persistent current also in non-ring geometries as racetracks confinements [426]. For stirring barrier strengths above a critical value, transitions from curved to straight racetrack parts are characterized by a circulation that results in an increase of one unit; in the transitions from straight to curved parts, instead the circulation results in a decrease by one unit. The analysis theoretically demonstrated that, playing with stirring barrier height, stir speed and racetrack geometry, quantized smooth atomic flows can be engineered on–demand in such systems. The interplay between curvature and atom-atom interactions in elliptic systems was explored in Refs. [427, 428].
@@
As was discussed in Sec. 4.1, the persistent flow is characterized by a hysteretic behavior as shown in Ref. [384] and shown in Fig. 12: when a rotating flow is made to decrease, the transition to lower winding numbers occurs at stirring frequencies smaller than the ones that were imparted to achieve the original rotation one started from [384]. The hysteresis results in a decrease as the barrier strength increases. While the quantitative understanding of such a phenomenon is still missing, there is a consensus in the community that hysteresis occurs as a result of the irreversibility encoded in vortex decay [392, 423, 429]. In Ref. [384], hysteresis among different branches is observed and proposed as a method to control an atomtronics device. In the mean-field regime, one can also investigate how hysteresis can be understood when following the ground state solution of a stirred BEC. In particular, in [429, 406], an exact analysis through GPE was shown to produce unstable solutions forming loops in the energy landscape of the system. As shown in Fig. 17, this leads to the facilitation of phase slips to access more stable states as the velocity of the barrier is increased above a critical velocity. Motivated by the implementations of the atomic counterpart of the dc-SQUID [138, 139], two moving weak links on the ring in opposite directions have been studied [387]. For such a configuration, vortex formation and phase slip dynamics explaining the current-phase relation was studied [425, 140]. Bosonic ring condensates with radii of a few hundred microns have been considered, therefore demonstrating the remarkable progress achieved in the coherent manipulation of condensates that can be seen with the naked eye.
@@ 5.1 BCS-BEC cross-over
A mixture of weakly attractive two-component Fermi gases forms pairs at opposite momenta around the Fermi sphere. This is the renowned Bardeen-Cooper-Schrieffer (BCS) state, which has been introduced to explain superconductivity. By tuning to stronger values the attraction among the fermions, the Cooper pairs modify onto smaller and finally close to point-like bosons, which can undergo Bose-Einstein condensation of such composite bosons. This is the celebrated BCS-BEC crossover, which plays an important role in different domains, ranging from nuclear to condensed matter physics. Ultracold atomic gases provide a unique platform to explore the BCS-BEC crossover as the interactions are tunable. This allows in particular to test different theoretical approximations. In the BCS regime, the correlation length of the bound pairs is large compared with the typical interparticle distance; in the BEC regime, instead, the pairs are tightly bound in the real space and the pair correlation length is much smaller than the distance between the particles. At intermediate values of the BCS-BEC crossover, one reaches a regime of infinitely attractive interactions, or unitary limit, where the gas properties are fixed just by the density and temperature of the gas.
@@ 5.2 Experimental results
Figure 18: (a) Absorption image of an equal spin mixture of lithium atoms in a ring-dimple trap, where there is a ring-shaped region of high chemical potential embedded in a broad halo of lower chemical potential. (b) Experimental radial column density. The black dotted line is the expected density profile for an ideal Fermi gas in the halo region. In region I, the system is 3D degenerate, II is quasi-2D degenerate, and III is quasi-2D thermal. The temperature of the system was inferred by fitting to the profile of the halo. Figure from Ref. [51].
@@
Faster interaction-driven ramps across the normal-to-superfluid phase transition were observed to result in spontaneous current formation, even for a non-rotating normal ring. This motivated the Dartmouth group to measure the rate of spontaneous current formation as a function of the quench rate, and compare the results against predictions obtained from Kibble-Zurek theory as shown in Ref. [360] and Fig. 19 (see also Sec. 4.1). In these experiments they drove thermal quenches as fast as 50 ms without substantial change in the peak density and chemical potential of the ring by using a hybrid quench protocol where the interaction strength and trap depth were ramped simultaneously. The trap geometry was a concentric double-ring, with one ring treated as a reference ring for interferometric detection of the current state in the other ring. For the fastest quench rates, the freeze-out time was reached before beyond-mean-field effects become significant and the spontaneous current formation rate matched well with predictions of mean-field Kibble-Zurek theory. For slower quenches, the rate of current production declined more rapidly. This was attributed to finite size effects, with the correlation length becoming comparable to the ring circumference for sufficiently slow quench rates.
@@
Figure 22: (a) After preparing an initial persistent current state with circulation $w_{0}$, the $n$ junctions are ramped up. The 3D density plots are isosurfaces obtained from 3D GPE numerical simulations of the experimental setup for the molecular BEC regime. If $n<n_{c}$ is below a critical value, the initial current is dissipated via the nucleation of vortices. Conversely, if $n\geq n_{c}$, the system remains stable with $w=w_{0}$ (lower right plot). (b) Examples of single-shot experimental in-situ images and interferograms obtained for $w_{0}=2$ and for the same number of junctions $n$ as in (a): n = 2 (unstable configuration), at $t=0$ (i), $t=1$ms (ii) and $t=7$ms (iii); and $n=4$ (stable configuration) for $t=0$ms (iv), $t=1$ms (v) and $t=20$ms (vi). In the case (iii), the circulation has decayed ($w(t)<w_{0}$) and the vortex emission is identified by the single spiral arm and the presence of a vortex. Figure from Ref. [261].
@@ 5.3 Theoretical Understanding
Still in its infancy, the theoretical study of persistent currents of ultracold matter in fermions is currently restricted to one-dimensional rings. At zero temperature and in the absence of interactions, the total energy of a spinless fermionic system is given by
@@
with $k_{n}=\frac{2\pi}{L}(n-\phi/\phi_{0})$ being the associated momentum of the eigenstates and $\{n\}$ being the set of quantum numbers corresponding to the energy levels occupied by the particles as shown in Fig. 23.(12)(12)(12)The set of quantum numbers $\{n\}$ is inherently related to the charge quantum number $I_{j}$ parameterizing the Bethe equations with $I_{j}=n$ or $I_{j}=n+\frac{1}{2}$ with $j=1,\ldots,N_{p}$. By construction, the sum of these charge quantum numbers is integer and is related to the angular momentum per particle as $\ell=N\frac{\sum_{j}I_{j}}{N_{p}}$. In contrast with bosons, only $N$ fermions can occupy a single-particle energy level due to the particle statistics. The resulting momentum distribution, centered compactly around zero to minimize the energy, is symmetric or asymmetric depending on whether the number of particles is $N_{p}=(2m+1)N$ or $N_{p}=(2m)N$ for integer $m$, i.e. $\frac{N_{p}}{N}$ is odd or even respectively, giving rise to the parity effect: a diamagnetic or paramagnetic current respectively –Fig. 23. Such a phenomenon is the straightforward generalization of the facets of the Leggett theorem discussed in Sec. 3.3 to $N$-component fermions [331, 88]. As the effective magnetic flux threading the system increases, the Fermi sphere is displaced so as to counteract it and minimize the energy as per Eq. (91[177, 90].
Figure 23: Schematic representation of the energy level occupations for non-interacting $N_{p}$ fermions with SU(2) symmetry and their behaviour with increasing angular momentum per particle $\ell$. To counteract the increase in $\ell$, the Fermi sphere displacement acquires a shift. Naturally, due to the Pauli exclusion, such a displacement for larger $N_{p}$ is mediated by going to higher $\ell$. Image reprinted from [90].
@@
Figure 24: Interference dynamics for weakly interacting SU($N$) fermions. (a) Homodyne protocol: Momentum distribution $\sum_{\{n\}}|J_{n}(\mathbf{k})|^{2}$ re-scaled by the number of components $N$ as defined in Eq. (76) with $n$ corresponding to level occupation of the particles. At angular momentum per particle $\ell=0$, the momentum distribution is peaked at the origin. Nevertheless, for $\ell>0$ there is a collapse, since the Fermi sphere is displaced as in Fig. 23 causing a characteristic hole to appear. The figure depicts three cases of equal number of particles per components: spinless, SU(2) and SU(3). Self-heterodyne protocol: Ring-centre interference $G_{R,C}$ for $N_{p}=6$ particles as a function of $N$ and $\ell$ at short time expansion $t=0.025$. Just as in its homodyne counterpart, the appearance of spirals necessitates that the Fermi sphere is displaced by half. For a fixed number of particles ($N_{p}=6$ in this case), the spirals emerge at lower $\ell$ (earlier value of the flux $\phi$) on increasing $N$. Additionally, dislocations (radially segmenting white lines) characterize the interferograms of different $N$. The colour bar is non-linear by setting $\mathrm{sgn}(G_{R,C})\sqrt{|G_{R,C}|}$. Images taken from [90].
@@
Figure 25: Persistent current $I(\phi)$ of SU($N$) symmetric fermions with repulsive interactions $U$ as a function of the synthetic gauge field $\phi$ for different filling fractions. (a) Incommensurate regime: Current exhibits a reduced periodicity when going to strong $U$. The quantities $\phi_{d}$ and $\phi_{s}$ correspond to the degeneracy point and flux at which spinons are generated respectively. (b) Mott regime: Persistent current as a function of $U$. Changes in its profile (magnitude and shape) capture the onset to the Mott phase transition. Figure taken from [88].
@@
The qualitative and quantitative changes of the persistent current make it a suitable probe to distinguish between the various bound states and to map out the system’s phase diagram [88, 92]. The current exhibits critical behaviour during a quantum phase transition despite its mesoscopic nature. For the case in question, the onset to the Mott phase transition in SU($N>2$) Hubbard models was demonstrated to be captured with the persistent current through finite-size scaling analysis [88]. Studies of Mott phase transitions with currents were also extended to multi-orbital SU($N$) Hubbard models to construct a conductivity-based phase diagram [96, 97, 95]. In contrast with the single-band model, it was shown that the current’s suppression with interaction is not always the case. Instead, it thrives in the deep insulating state, and in certain regions known as Hund’s metal region, it is also enhanced as the hopping processes across the different orbitals facilitate the fermions’ motion even at strong interactions.
@@
Figure 28: Homodyne protocol for strongly interacting SU($N$) fermionic systems. (a) Momentum distribution cross-section $n(k_{x},0)$ for $N_{p}=3$ particles with large repulsive interactions $U=10,000$ as a function of the effective flux $\phi$. The initially peaked momentum distribution collapses at $k_{x}=0$ once the additional threshold constraint $(N_{p}-1)/2N_{p}$ enforced by the fractionalization is cleared. Insets correspond to the full momentum distribution. (b) Variance of the width of the momentum distribution $\sigma_{n_{k}}(\phi)$ for strongly attractive fermions. The appearance of discrete steps per flux quantum $\phi_{0}$ corresponds to the number of particles composing the bound state and ($N$ in the case shown), in turn, reflects the angular momentum fractionalization. Figure taken from [90].
In the presence of finite temperature, changes in the persistent current periodicity stem from the interplay between interactions and thermal fluctuations. Employing a thermodynamic Bethe ansatz approach for repulsively interacting two-component fermions, it was demonstrated that apart from having its sawtooth shape smoothened out, the persistent current’s amplitude is also exponentially reduced with the decay rate dependent on the system’s parameters [437]. The suppression of the current could be attributed to temperature activated transitions between states with different angular momenta. Alternatively, the persistent current contribution from the populated levels are equal in magnitude but with different directionalities thereby cancelling each other out. A remarkable effect of finite temperature is that the single-particle frequency of the persistent current is re-instated, imitating, in a way, a reduction in the interaction strength – Fig. 27. This ‘de-fractionalization’ phenomenon was also demonstrated to hold for $N$-component fermions with attractive interactions [89]. The mechanism behind this is that the contributing excited states have different frequencies that modify the current’s periodicity, with the general trend being that the system becomes more robust to thermal fluctuations with increasing interactions. Nevertheless, for strong attractive interactions the system is more susceptible to temperature. It turns out that the bound states sub-band, which is the only contributor to the current in this regime, flattens out due to a decrease in the level spacing between the states constituting it. Consequently, at small temperatures the system exists as a ‘gas of bound states’ altering the persistent current periodicity in the same manner discussed previously.
@@
To round things off, we consider the read-out of persistent currents in the strongly interacting regimes through interferometric means.(13)(13)(13)From the theoretical standpoint, the calculation of interference patterns for strongly interacting systems is limited to small particle numbers and system sizes. Unlike the free fermion regimes, there is no exact expression for the correlators in this regime and these systems are intractable by numerical methods such as exact diagonalization and DMRG for large particle numbers and system sizes. Nevertheless for the strongly repulsive interacting regime, it was shown that by employing a Bethe ansatz scheme to split the SU($N$) Hubbard model into the spinless fermionic and SU($N$) Heisenberg models, interference patterns for systems well beyond the current state-of-the-art numerical methods can be obtained [91]. In the homodyne protocol, the fractionalization in the system generates a momentum distribution depression, i.e. a non-zero local minimum in the origin, instead of the characteristic hole. Furthermore, the appearance of this depression occurs at larger flux values compared to the zero and weak interaction picture. On top of the Fermi sphere effect explained previously, there is an additional delay of $\frac{N_{p}-1}{2N_{p}}$ ($\frac{N-1}{2N}$) in strongly repulsive (attractive) regimes, which is a signature of the fractionalization –Fig. 28(a). The visibility of the depression for attracting fermions is significantly diminished due to the reduced coherence associated with the bound state formation. Nevertheless, information can be acquired indirectly from the variance of the width of the momentum distribution, with the appearance of discrete steps whose number in a given flux quantum correspond to the angular momentum fractionalization –Fig. 28(b). For the self-heterodyne protocol, the interferograms also embody the features of the fractionalized with the emergence of the spiral corresponding with the depression in the momentum distribution. Aside from being able to track the angular momentum quantization by spiral’s specific delay, it can also be monitored through the change in the number and orientation of the dislocations. At variance with free fermions, the dislocations in the strong interaction regime are flux dependent being modified by the presence of spin excitations giving a visual cue about the angular momentum fractionalization –Fig. 29. In the case of attractive interactions, self-heterodyne interference patterns are not a good figure of merit for SU($N>2$) systems since the features of $N$-body bound are not captured by the density-density correlators used to generate these interferograms.